What is your maximum possible profit if you long one Apple, Inc. July 74 put contract with a premium of $2.59?
Answer #1

We are long the put option, that gives us the right to sell the underlying stock at the exercise price of 74. Now whenever the price goes below 74, we will exercise the option. So lowest price that any stock can go is 0 so if that is the case then the payoff is 74-0 = 74, however, we have paid the premium of 2.59 to buy the long option, so reducing that from the payoff, the profit = 74-2.59 = 71.41